OpenStudy (anonymous):

A government has 5.7 x 10 to the 7th power dollars invested in a pension fund for retired employees. It expects the investment to double in value every 8 years. What is the value of the investment after 8 years, 16 years, and 24 years? Write your response in scientific notation.

OpenStudy (callisto):

After 8 year , amount = \[2\times 5.7\times10^{7} = 1.14\times10^8\] after 16 years amount = \[2\times 1.14\times10^{8} = 2.28\times10^8\] after 24years amount = \[2\times 2.28\times10^{8} = 4.56\times10^8\]